# Cohomology ring of non-compact abelian arrangements

### v1

@@ 3. Central and unimodular arrangements
The horizontal maps in the left-hand square are induced by the inclusion of the complement of $\mathcal{A}$ in the complement of $\mathcal{B}$, in $\mathbb{R}^{n}$ and $G^{n}$ respectively. Since $M^{\mathbb{R}}(\mathcal{B})\cong(\mathbb{R}\setminus 0)^{n}$ and $M^{a,b}(\mathcal{B})\cong(G\setminus e)^{n}$, the horizontal maps in the right-hand square are given by the Künneth isomorphisms. Note that the Künneth isomorphism $k^{a,b}$ depends on the sign of the basis $B$.
For the sake of notation we assume that $I$ are the first elements of $B$, the general case differs by the sign $(-1)^{d\ell(I,B\setminus I)}$ given by the reordering of $\omega_{i}$ and $\psi_{b}$ for $i\in I$ and $b\in B\setminus I$. We have
@@
where in the last equality we use Remark 3.6. If $b>1$, $i_{*}(1)=0$ and hence ${i_{\mathcal{A}}}_{*}(w^{+}_{I})=0$ if $\operatorname{rk}(I)\neq n$. Otherwise, $I$ is a basis and
@@
Remark 3.8.
Note that, also in the case $b=1$, the image of $i_{*}$ does not depend on the choice of the (unimodular) basis $B$. In fact, if $B^{\prime}\subseteq E$ is another (unimodular) basis containing $I$, $\operatorname{sgn}(B)^{d}\psi_{B\setminus I}\omega_{I}=\operatorname{sgn}(B^{\prime})^{d}\psi_{B^{\prime}\setminus I}\omega_{I}$ by using the linear relations among the classes $\psi$ coming from the linear dependencies among the characters of $B$ and $B^{\prime}$.
@@
A subset $A\subseteq\mathcal{A}$ is a generalized circuit if $\operatorname{rk}(A)=|A|-1$. We denote by $C(A)$ the unique circuit contained in $A$.
Definition 3.10.
Let $C\subseteq E$ be a circuit and let $B\supset C$ such that $B\setminus{\{i\}}$ is a basis for any $i\in C$. We denote by $c_{i}$ the sign of the determinant of the matrix whose columns are the elements in $B\setminus{\{i\}}$.
Remark 3.11.
Up to a global sign, the $c_{i}$’s do not depend on the choice of $B$. The collection of these signs defines a chirotope (or equivalently an oriented matroid), cf [${}^{+}$99].
@@
    $\sum_{\begin{subarray}{c}K\subseteq C^{-}\\ K\neq\emptyset\end{subarray}}(-1)^{|K|}c_{i_{K}}^{d}\eta_{C\setminus K,K\setminus i_{K}}-\sum_{\begin{subarray}{c}K\subseteq C^{+}\\ K\neq\emptyset\end{subarray}}(-1)^{|K|}c_{i_{K}}^{d}\eta_{C\setminus K,K\setminus i_{K}}=0,$ (9)
@@
We consider the relation (8) in $H^{0}(M^{\mathbb{R}}(C))$ and we apply Lemma 3.7 to the essentialization of the arrangement $C$. For any independent $J\subsetneq C^{+}$ (resp. $J\subsetneq C^{-}$), the choice of $i_{K}\in K=C^{+}\setminus J$ (resp. $i_{K}\in K=C^{-}\setminus J$) corresponds to complete $C^{+}\bigsqcup J=C\setminus K$ to a basis $C\setminus i_{K}$ of the sublattice generated by the elements of $C$. By Lemma 3.7 and Remark 3.6 we have:
@@
Example 3.14.
{tikzpicture}{tikzpicture}
Figure 2. A picture of the arrangement in Example 3.14 for $(a,b)=(1,0)$ on the left, and $(a,b)=(0,1)$ on the right.
@@
(see Figure 2). This is an arrangement of type $A_{3}$, hence $M(\mathcal{A})$ is the configuration space of $3$ points in $G$ (c.f. Section 6). The subspace $H^{*}(G^{2})\subset H^{*}(M(\mathcal{A}))$ is generated by the classes $\psi_{1}^{j},\psi_{2}^{j}$ and $\psi_{3}^{j}=\psi_{1}^{j}+\psi_{2}^{j}$, with $j=1,\cdots,a$. Furthermore, we have defined the classes $\omega_{i}\in H^{a+b-1}(M(\mathcal{A}))$ for $i=1,2,3$. The circuit $C=C^{+}\sqcup C^{-}=\{1,2\}\sqcup\{3\}$ gives the following relations
@@
4. Non-central arrangements
Now, we deal with the case of non-central totally unimodular arrangements. Firstly, we prove a variation of the Brieskorn lemma, a deletion-restriction short exact sequence, and an identity between Poincaré and characteristic polynomial. We will use these results in Section 5 for unimodular covers.
Given $H\in\mathcal{A}$ a subvariety of the arrangement, we denote by $\mathcal{A}^{\prime}:=\mathcal{A}\setminus\{H\}$ the deletion of $H$ and $\mathcal{A}^{\prime\prime}$ the restriction to $H$, i.e. the arrangement of connected components of $K\cap H$ for $K\in\mathcal{A}^{\prime}$. For any layer $p\in\mathcal{L}(\mathcal{A})$ we consider the local arrangement $\mathcal{A}_{p}:=\{H\in\mathcal{A}\mid p\subseteq H\}$ at $p$ and let $g_{p}\colon M(\mathcal{A})\to M(\mathcal{A}_{p})$ be the inclusion of the complements.
@@ 4. Non-central arrangements
In the central case the theorem follows from [LTY21, Theorem 7.6]. We prove the claims by induction on the cardinality of the arrangement $\mathcal{A}$. The base case is a central arrangement and the results hold. The assumption that $\mathcal{A}$ is not central implies $\lvert\mathcal{A}_{p}\rvert<\lvert\mathcal{A}\rvert$ for any $p\in\mathcal{L}^{r}$. Consider $p\in\mathcal{L}(\mathcal{A}^{\prime\prime})$ and let $\tilde{p}\in\mathcal{L}(\mathcal{A}^{\prime})$ be the unique minimal layer containing $p$ (if $H$ is not a coloop then $p=\tilde{p}$), the diagram
{tikzcd}
is a pullback diagram and so $\iota_{*}\circ g^{\prime\prime*}_{p}=g^{\prime*}_{\tilde{p}}\circ\iota_{p*}$. Since the bundary map of the long exact sequence and the Thom isomorphism are functorial, the following diagram commutes:
{tikzcd}
@@
{tikzcd}
because $\mathcal{A}^{\prime}_{p}=\mathcal{A}_{p}$. We consider the direct sum of previous sequences for all $p\in\mathcal{L}(\mathcal{A})$. Since $\lvert\mathcal{A}^{\prime\prime}\rvert<\lvert\mathcal{A}\rvert$, the map $\oplus_{p\in\mathcal{L}^{r}}g_{p}^{\prime\prime*}$ is surjective and $\iota_{p*}=0$ by inductive step:
{tikzcd}
@@
$0\to H^{*}(M(\mathcal{A}^{\prime}))\xrightarrow{j^{*}}H^{*}(M(\mathcal{A}))\xrightarrow{\operatorname{res}}H^{*-(a+b-1)}(M(\mathcal{A}^{\prime\prime}))\to 0$
@@
For any abelian arrangement $\mathcal{A}$, the Poincaré polynomial only depends on the characteristic polynomial of the arrangement:
@@
The identity follows by induction using the exact sequence in Theorem 4.1. ∎
@@
Example 4.4.
{tikzpicture}{tikzpicture}
Figure 3. A picture of the non central arrangement in Example 4.4 when $(a,b)=(1,0)$ on the left, and $(a,b)=(0,1)$ on the right.
Let us consider the arrangement obtained from the arrangement in the example 3.14 by adding the two hyperplanes $H_{2^{\prime}}=\{\chi_{2}^{-1}(-e)\}$ and $H_{3^{\prime}}=\{(\chi_{1}+\chi_{2})^{-1}(-e)\}$ as shown in Figure 3.
@@
When $b=1$, the first circuit leads to the same relation eq. (12) of Example 3.14, while the second one to the new relation:
@@ 5. Non-unimodular case
The poset of layers of a non-unimodular arrangement $\mathcal{A}$ is no longer a geometric lattice, due to multiple connected components of the intersections. Indeed, the combinatoric data of $\mathcal{A}$ are encoded in a generalization of a matroid, basically a matroid together with a multiplicity function, known as arithmetic matroid and first introduced by L. Moci and M. D’Adderio in [DM13]. We deal with the non-unimodular case adapting the argument in [${}^{+}$20], by constructing an unimodular covering for an abelian arrangement $\mathcal{A}$.
@@
Let $\Lambda$ be a lattice and $\mathcal{A}$ an arrangement in $G^{\Lambda}$, define the multiplicity
$m(A)=[\Lambda^{A}:\Lambda_{A}]$
where $\Lambda_{A}:=\langle A\rangle\subseteq\Lambda$ and $\Lambda^{A}$ is the radical of $\Lambda_{A}$ in $\Lambda$, i.e. $\Lambda^{A}=(\mathbb{Q}\otimes_{\mathbb{Z}}\Lambda_{A})\cap\Lambda$. This number $m(A)$ is the cardinality of the torsion subgroup of $\Lambda/\Lambda_{A}$ and the multiplicity function of the matroid associated to $\mathcal{A}$.
Suppose $\mathcal{A}=\{\chi_{0},\cdots,\chi_{r}\}$ consists of one generalized circuit $X=\{0,\cdots,r\}=C\sqcup F$, with $n=\operatorname{rk}(C)$. Let
@@
and let $\Lambda^{\prime}$ be the lattice in $\Lambda\otimes\mathbb{Q}$ generated by the characters $\frac{\chi_{i}}{a_{i}}$.
@@
The lattice Λsuperscriptnormal-$\Lambda^{\prime}$ contains $\Lambda$ and the inclusion induces a covering $\pi\colon G^{\Lambda^{\prime}}\rightarrow G^{\Lambda}$ of degree $m(C)^{a}m(X)^{a(r-1)}\prod_{i\in C}m(C\setminus i)^{a(n-1)}$.
@@
The proof of the first statement is analogous to that of [${}^{+}$20, Lemma 6.4]. Let $i\in C$, as in [${}^{+}$20, Lemma 6.5] we compute the index of $\Lambda$ in $\Lambda^{\prime}$
@@
Let $\mathcal{A}_{U}$ be the arrangement in $U:=G^{\Lambda^{\prime}}$ given by the set of connected components of the preimages of the subvarieties in $\mathcal{A}$
@@
Let $A\subseteq E$, $W$ a connected component of $\bigcap_{i\in A}H_{i}$ and $p\in W$. Since $\pi^{*}$ is injective, we define $\omega_{W,A}$ as the unique class in $H^{*}(M(\mathcal{A}))$ such that
$\pi^{*}(\omega_{W,A})=\frac{1}{|L\cap\pi^{-1}(p)|}\sum_{q\in\pi^{-1}(p)}\omega_{A}^{U}(q)=\frac{1}{|L\cap\pi^{-1}(p)|}\sum_{\tilde{L}\text{ c.c. of }\pi^{-1}(W)}\omega_{A}^{U}(q_{\tilde{L}})$ (18)
@@
Let $A,B\subseteq E$ such that $A\cup B$ is a dependent set and let $W$ be a connected component of $\bigcap_{i\in A}H_{i}$. Then, for any $j=1,\cdots,a$,
$\omega_{W,A}\psi^{j}_{B}=0.$ (19)
@@
By definition, it is enough to prove that, for any connected component $L$ of $\pi^{-1}(W)$, $\omega^{U}_{A}(q_{L})\psi_{B}^{j,U}=0$ in $H^{*}(M(\mathcal{A}_{U}))$. This follows from $(\psi_{i}^{j,U})^{2}=0$, $\omega_{i}^{U}(q_{L})\psi_{i}^{j,U}=0$ and from the linear dependences between $\{\chi_{i},i\in A\cup B\}$. ∎
@@
Let $X$ be a generalized circuit and $A,B\subseteq X$ such that $A\sqcup B$ is a maximal independent set. Let $W$ be a connected component of $\bigcap_{i\in A}H_{i}$ and $p\in W$. Then,
@@
The computation of $|L\cap\pi^{-1}(p)|$ is analogous to that in Lemma 6.8 in [${}^{+}$20]. The cardinality of the preimage of $p$ is the degree of the covering. The number of connected components of $\pi^{-1}(W)$ is
@@
In this last computation the final equality comes from property $(3)$ of arithmetic matroids, see [DM13]. ∎
@@
Let $X=C\sqcup F$ be a generalized oriented circuit, $Y$ be a connected component of $\bigcap_{i\in X}H_{i}$, then
@@
We apply Lemma 3.12 to the circuit $C$ in the abelian variety $U=G^{\Lambda^{\prime}}$. Let $p\in Y$ be any point. If $b=1$, we have:

### v2

@@ 3. Central and unimodular arrangements
The horizontal maps in the left-hand square are induced by the inclusion of the complement of $\mathcal{A}$ in the complement of $\mathcal{B}$, in $\mathbb{R}^{r}$ and $G^{r}$ respectively. Since $M^{\mathbb{R}}(\mathcal{B})\cong(\mathbb{R}\setminus\{0\})^{r}$ and $M^{a,b}(\mathcal{B})\cong(G\setminus\{e\})^{r}$, the horizontal maps in the right-hand square are given by the Künneth isomorphisms. Note that the Künneth isomorphism $k^{a,b}$ depends on the sign of the basis $B$.
For the sake of notation we assume that $I$ are the first elements of $B$, the general case differs by the sign $(-1)^{d\ell(I,B\setminus I)}$ (where $\ell(A,B)$ is introduced in Definition 3.2) given by the reordering of $\omega_{i}$ and $\psi_{b}$ for $i\in I$ and $b\in B\setminus I$. We have
@@
where in the last equality we use Remark 2. Otherwise $b>1$, $i_{*}(1)=0$ and hence ${i_{\mathcal{A}}}_{*}(w^{+}_{I})=0$ if $\operatorname{rk}(I)\neq r$. If $\operatorname{rk}(I)=r$, then
@@
Remark 3.
Note that, also in the case $b=1$, the image of $i_{*}$ does not depend on the choice of the (unimodular) basis $B$. In fact, if $B^{\prime}\subseteq E$ is another (unimodular) basis containing $I$, $\operatorname{sgn}(B)^{d}\psi_{B\setminus I}\omega_{I}=\operatorname{sgn}(B^{\prime})^{d}\psi_{B^{\prime}\setminus I}\omega_{I}$ by using the linear relations among the classes $\psi$.
@@
Let $C\subseteq E$ be a circuit and $B\supseteq C$ such that $B\setminus{\{i\}}$ is a basis for any $i\in C$. We denote by $c_{i}$ the sign of the determinant of the matrix whose columns are the elements in $B\setminus{\{i\}}$.
Up to a global sign, the $c_{i}$’s do not depend on the choice of $B$.
@@
    $\sum_{\begin{subarray}{c}K\subseteq C^{-}\\ K\neq\emptyset\end{subarray}}(-1)^{|K|}c_{i_{K}}^{d}\eta_{C\setminus K,K\setminus i_{K}}-\sum_{\begin{subarray}{c}K\subseteq C^{+}\\ K\neq\emptyset\end{subarray}}(-1)^{|K|}c_{i_{K}}^{d}\eta_{C\setminus K,K\setminus i_{K}}=0$ (5)
@@
Consider the relation (4) in $H^{0}(M^{\mathbb{R}}(C))$ and we apply Lemma 3.4 to the essentialization of the arrangement $C$. For any independent $J\subsetneq C^{+}$ (resp. $J\subsetneq C^{-}$), the choice of $i_{K}\in K=C^{+}\setminus J$ (resp. $i_{K}\in K=C^{-}\setminus J$) corresponds to complete $C^{+}\bigsqcup J=C\setminus K$ to a basis $C\setminus i_{K}$ of the sublattice generated by the elements of $C$. By Lemma 3.4 and Remark 2 we have:
@@
Example 1.
[…]
@@
(see Figure 2). This is an arrangement of type $A_{3}$, hence $M(\mathcal{A})$ is the configuration space of 333 points in $G$ (c.f. Section 7). The subspace $H^{*}(G^{2})\subset H^{*}(M(\mathcal{A}))$ is generated by the classes $\psi_{1}^{j},\psi_{2}^{j}$ and $\psi_{3}^{j}=\psi_{1}^{j}+\psi_{2}^{j}$, with $j=1,\cdots,a$. Furthermore, we have defined the classes $\omega_{i}\in H^{a+b-1}(M(\mathcal{A}))$ for $i=1,2,3$. The circuit $C=C^{+}\sqcup C^{-}=\{1,2\}\sqcup\{3\}$ gives the following relations
@@
4. The long exact sequence of the pair
Now, we deal with the case of general arrangements, possibly non-central and non-unimodular. Firstly, we prove a variation of the Brieskorn Lemma, a deletion-restriction short exact sequence, and an identity between Poincaré and characteristic polynomial. We will apply these results in Section 5 to unimodular covers.
Given $H\in\mathcal{A}$ a subvariety of the arrangement, we denote by $\mathcal{A}^{\prime}:=\mathcal{A}\setminus\{H\}$ the deletion of $H$ and $\mathcal{A}^{\prime\prime}$ the restriction to $H$, i.e. the arrangement of connected components of $K\cap H$ for $K\in\mathcal{A}^{\prime}$. The complement $M(\mathcal{A}^{\prime})=M(\mathcal{A})\sqcup M(\mathcal{A}^{\prime\prime})$ is a disjoint union of an open and a closed subset. The relative cohomology $H^{*}(M(\mathcal{A}^{\prime}),M(\mathcal{A}))$ is isomorphic to $H^{*-a-b}(M(\mathcal{A}^{\prime\prime}))$ via the Thom isomorphism and excision theorem. Therefore, the forgetful map of the long exact sequence of the pair $(M(\mathcal{A}^{\prime}),M(\mathcal{A}))$ can be identified through the aforementioned isomorphism with the pushforward $\iota_{*}$, where $\iota\colon M(\mathcal{A}^{\prime\prime})\to M(\mathcal{A}^{\prime})$ is the closed immersion (see property (P4)).
For any layer $p\in\mathcal{L}(\mathcal{A})$ we consider the local arrangement $\mathcal{A}_{p}:=\{H\in\mathcal{A}\mid p\subseteq H\}$ at $p$ and let $g_{p}\colon M(\mathcal{A})\to M(\mathcal{A}_{p})$ be the inclusion of the complements.
@@ 4. The long exact sequence of the pair
In the central case the theorem follows from [LTY21, Theorem 7.7]. We prove the claims by induction on the rank and the cardinality of the arrangement $\mathcal{A}$. The base case is a central arrangement and the results hold. The assumption that $\mathcal{A}$ is not central implies $\lvert\mathcal{A}_{p}\rvert<\lvert\mathcal{A}\rvert$ for any $p\in\mathcal{L}^{r}$. Consider $p\in\mathcal{L}(\mathcal{A}^{\prime\prime})$ and let $\tilde{p}\in\mathcal{L}(\mathcal{A}^{\prime})$ be the unique minimal layer containing $p$ (if $H$ is not a coloop then $p=\tilde{p}$), the diagram
${M(\mathcal{A}^{\prime\prime})}$${M(\mathcal{A}^{\prime})}$${M(\mathcal{A}^{\prime\prime}_{p})}$${M(\mathcal{A}^{\prime}_{\tilde{p}})}$$\scriptstyle{\iota}$$\scriptstyle{g^{\prime\prime}_{p}}$$\scriptstyle{g^{\prime}_{\tilde{p}}}$$\scriptstyle{\iota_{p}}$
is a pullback diagram and so $\iota_{*}\circ g^{\prime\prime*}_{p}=g^{\prime*}_{\tilde{p}}\circ\iota_{p*}$ by property (P3). Since the long exact sequence of the pair and the Thom isomorphism are functorial, the following diagram commutes:
${H^{k-d-1}(M(\mathcal{A}_{p}^{\prime\prime}))}$${H^{k}(M(\mathcal{A}^{\prime}_{\tilde{p}}))}$${H^{k}(M(\mathcal{A}_{p}))}$${H^{k-d}(M(\mathcal{A}_{p}^{\prime\prime}))}$${H^{k-d-1}(M(\mathcal{A}^{\prime\prime}))}$${H^{k}(M(\mathcal{A}^{\prime}))}$${H^{k}(M(\mathcal{A}))}$${H^{k-d}(M(\mathcal{A}^{\prime\prime}))}$$\scriptstyle{\iota_{p}^{*}}$$\scriptstyle{g_{p}^{\prime\prime*}}$$\scriptstyle{j_{p}^{*}}$$\scriptstyle{g^{\prime*}_{\tilde{p}}}$$\scriptstyle{\delta_{p}}$$\scriptstyle{g^{*}_{p}}$$\scriptstyle{g_{p}^{\prime\prime*}}$$\scriptstyle{\iota_{*}}$$\scriptstyle{j^{*}}$$\scriptstyle{\delta}$
@@
[…]
because $\mathcal{A}^{\prime}_{p}=\mathcal{A}_{p}$. We consider the direct sum of previous sequences for all $p\in\mathcal{L}(\mathcal{A})$. Since $\operatorname{rk}(\mathcal{A}^{\prime\prime})<\operatorname{rk}(\mathcal{A})$, the map $\oplus_{p\in\mathcal{L}^{r}}g_{p}^{\prime\prime*}$ is surjective and $\iota_{p*}=0$ by inductive step:
[…]
@@
$0\to H^{*}(M(\mathcal{A}^{\prime}))\xrightarrow{j^{*}}H^{*}(M(\mathcal{A}))\xrightarrow{\operatorname{res}}H^{*-d}(M(\mathcal{A}^{\prime\prime}))\to 0$
The characteristic polynomial of an abelian arrangement is
$\chi_{\mathcal{A}}(t):=\sum_{W\in\mathcal{L}(\mathcal{A})}\mu(\hat{0},W)t^{r-\operatorname{rk}(W)},$
where $\mu$ is the Möbius function of the poset of layers $\mathcal{L}(\mathcal{A})$.
@@
For any abelian arrangement $\mathcal{A}$ in a non-compact group $G$, the Poincaré polynomial only depends on the characteristic polynomial of the arrangement:
@@
The identity follows by deletion-contraction recurrence, using the exact sequence in Corollary 4.2. ∎
@@
Corollary 4.4.
Let $C$ be a unimodular, possibly non-central, circuit of $\mathcal{A}$. Then
$\omega_{C}=0.$
Proof.
Assume that $C=\{1,2,\dots,n\}$ and consider the subvariety $H_{0}$ obtained as the unique translated of $H_{1}$ containing $\cap_{i\in C\setminus\{1\}}H_{i}$. The set $D:=(C\cup\{0\})\setminus\{1\}$ is a central circuit and we can multiply the relations of Lemma 3.6 by $\omega_{1}$ obtaining
$\sum_{\begin{subarray}{c}K\subseteq D^{-}\\ K\neq\emptyset\end{subarray}}(-1)^{|K|}c_{i_{K}}^{d}\omega_{1}\eta_{D\setminus K,K\setminus i_{K}}-\sum_{\begin{subarray}{c}K\subseteq D^{+}\\ K\neq\emptyset\end{subarray}}(-1)^{|K|}c_{i_{K}}^{d}\omega_{1}\eta_{D\setminus K,K\setminus i_{K}}=0.$
The elements 0,10,10 , 1 are parallel, so $\omega_{0}\omega_{1}=0$ and $\omega_{1}\psi_{0}=0$, see also Remark 2. In particular, all the terms in the above sum are zero unless the one with $K=\{0\}$, hence the equation reduces to $\omega_{C}=0$ in the cohomology ring of the arrangement $\mathcal{A}\cup\{H_{0}\}$. The claimed equality follows from the injectivity of the map $H^{*}(M(\mathcal{A}))\to H^{*}(M(\mathcal{A}\cup\{H_{0}\}))$ by Corollary 4.2. ∎
Example 2.
[…]
Let us consider the arrangement obtained from the arrangement in the Example 1 by adding the two hyperplanes $H_{2^{\prime}}=\{\chi_{2}^{-1}(-e)\}$ and $H_{3^{\prime}}=\{(\chi_{1}+\chi_{2})^{-1}(-e)\}$ as shown in Figure 3.
@@
When $b=1$, the first circuit leads to the same relation eq. (8) of Example 1, while the second one to the new relation:
@@ 5. Non-unimodular case
In this section we prove our main result on the cohomology of the complement of abelian arrangements, without any assumption of centrality and unimodularity. We adapt the argument in [CDD+20] constructing an separating cover for an abelian arrangement $\mathcal{A}$.
@@
A separating cover of $A\subseteq E$ is a cover $\pi\colon\operatorname{Hom}(\Gamma,G)\to\operatorname{Hom}(\Lambda,G)$ such that the connected components of $\pi^{-1}(H_{i})$ for $i\in A$ form a unimodular, possibly non-central, arrangement.
A covering is induced from an extension $\Gamma\supset\Lambda$, for each character $\chi\in\Lambda$ let $a\in\mathbb{N}$ be biggest integer such that $\frac{\chi}{a}\in\Gamma$. The cover separates $A$ if and only if the characters $\frac{\chi_{i}}{a_{i}}$ generate a split subgroup of $\Gamma$. In general, there is no separating cover for $A$, but if $A$ has nullity at most one then it exists.
Indeed, let $X\subseteq E$ of nullity 111, hence $X=C\sqcup F$ contains a unique circuit $C$ and set $n=\operatorname{rk}(C)$. Let
@@
and let $\Lambda^{\prime}$ be the lattice in $\Lambda\otimes_{\mathbb{Z}}\mathbb{Q}$ generated by the characters $\frac{\chi_{i}}{a_{i}}$. This cover is separating for $X$ of nullity 111.
@@
The lattice $\Lambda^{\prime}$ contains $\Lambda$ and the inclusion induces a covering $\pi\colon\operatorname{Hom}(\Lambda^{\prime},G)\rightarrow\operatorname{Hom}(\Lambda,G)$ of degree $m(C)^{a}m(X)^{a(r-1)}\prod_{i\in C}m(C\setminus i)^{a(n-1)}$.
@@
The proof of the first statement is analogous to that of [CDD+20, Lemma 6.4]. Let $i\in C$, as in [CDD+20, Lemma 6.5] we compute the index of $\Lambda$ in $\Lambda^{\prime}$
@@
Let $\mathcal{A}_{U}$ be the arrangement in $U:=\operatorname{Hom}(\Lambda^{\prime},G)$ given by the set of connected components of the preimages of the subvarieties in $\mathcal{A}$
@@
Let $A\subseteq E$ be a independent subset, $W$ a connected component of $\bigcap_{i\in A}H_{i}$ and $p\in W$. Since $\pi^{*}$ is injective, we define $\omega_{W,A}$ as the unique class in $H^{*}(M(\mathcal{A});\mathbb{Z})$ such that
$\pi^{*}(\omega_{W,A})=\frac{1}{|L\cap\pi^{-1}(p)|}\sum_{q\in\pi^{-1}(p)}\omega_{A}^{U}(q)=\sum_{\tilde{L}\text{ c.c. of }\pi^{-1}(W)}\omega_{A}^{U}(q_{\tilde{L}})$ (14)
@@
The classes $\omega_{W,A}$ do not depend on the choice of the cover $\pi\colon U\to G^{r}$ as in [CDD+20, Theorem 5.4].
@@
Let $A$ be an independent set and $W$ a connected component of $\bigcap_{i\in A}H_{i}$. Then, for every $z\in\ker\left(H^{*}(G^{r};\mathbb{Z})\to H^{*}(W;\mathbb{Z})\right)$
$\omega_{W,A}z=0.$
In particular, if $B\subseteq E$ is dependent on $A$ then for any $j=1,\cdots,a$,
$\omega_{W,A}\psi^{j}_{B}=0.$
@@
If $A$ is a unimodular subset then $\omega_{W,A}=\omega_{a_{1}}\omega_{a_{2}}\dots\omega_{a_{k}}$ and $\omega_{W,A}\psi_{i}^{j}=0$ for every $i\in A$ and any $j\leq a$. Such elements $\psi_{i}^{j}$ generates $\ker\left(H^{*}(G^{r};\mathbb{Z})\to H^{*}(W;\mathbb{Z})\right)$ as an ideal and hence $\omega_{W,A}z=0$.
In the non unimodular case, we consider the separating cover $\mathcal{A}_{U}$ and let $L$ be a connected component of $\pi^{-1}(W)$. Since $\pi^{*}$ is injective, it is enough to prove that $\omega^{U}_{A}(q_{L})z^{U}=0$ for any $z^{U}\in\ker\left(H^{*}(U;\mathbb{Z})\to H^{*}(L;\mathbb{Z})\right)$, because $\pi^{*}(\omega_{W,A})$ is a linear combination of $\omega^{U}_{A}(q_{L})$. The arrangement $\mathcal{A}_{U}$ is unimodular and the result follows as in the previous case.
The last statement follows from the fact that $\psi_{B}^{j}\in\ker\left(H^{*}(G^{r};\mathbb{Z})\to H^{*}(W;\mathbb{Z})\right)$ if (and only if) the set $B$ is dependent on $A$. ∎
@@
Let $X$ be a set of nullity 111 and $A,B\subseteq X$ such that $A\sqcup B$ is a maximal independent set. Let $W$ be a connected component of $\bigcap_{i\in A}H_{i}$ and $p\in W$. Then,
@@
The following computation of $|L\cap\pi^{-1}(p)|$ is analogous to that in [CDD+20, Lemma 6.10]. The cardinality of the inverse image of $p$ is the degree of the covering. The number of connected components of $\pi^{-1}(W)$ is
@@
In this last computation the final equality comes from property (AM3) of arithmetic matroids. ∎
@@
We have
Proof.
We start with the case of $A\sqcup A^{\prime}$ of nullity at most one. Consider any separating cover $\pi\colon U\to G^{r}$ of $A\sqcup A^{\prime}$, we have
$\pi^{*}(\omega_{W,A}\omega_{W^{\prime},A^{\prime}})$ $=\sum_{L\in\pi_{0}(\pi^{-1}(W))}\omega_{A}^{U}(q_{L})\sum_{L^{\prime}\in\pi_{0}(\pi^{-1}(W^{\prime}))}\omega_{A^{\prime}}^{U}(q_{L^{\prime}})$
$=\sum_{K\in\pi_{0}(\pi^{-1}(W\cap W^{\prime}))}\omega_{A}^{U}(q_{K})\omega_{A^{\prime}}^{U}(q_{K})$
$=(-1)^{d\ell(A,A^{\prime})}\sum_{L\in\pi_{0}(W\cap W^{\prime})}\sum_{K\in\pi_{0}(\pi^{-1}(L))}\omega_{A\sqcup A^{\prime}}^{U}(q_{K})$
$=(-1)^{d\ell(A,A^{\prime})}\sum_{L\in\pi_{0}(W\cap W^{\prime})}\pi^{*}(\omega_{L,A\sqcup A^{\prime}})$
if $A\cap A^{\prime}=\emptyset$, otherwise in the second line above we have $\omega_{A}^{U}(q_{K})\omega_{A^{\prime}}^{U}(q_{K})=0$ because $(\omega_{i}^{U}(q_{K}))^{2}=0$. By injectivity of $\pi_{*}$ and Corollary 4.4 the claimed equality follows.
Now consider arbitrary independent sets $A$ and $A^{\prime}$ and fix a subset $A\subset X\subseteq A\cap A^{\prime}$ of nullity one. Consider a separating cover $f\colon U\to G^{r}$ for $A^{\prime}$ and a separating cover $g\colon V\to U$ for $f^{-1}(X)$. Notice that $f\circ g$ is a separating cover for $X$. We have
$f^{*}(\omega_{W^{\prime},A^{\prime}})=\sum_{L^{\prime}\in\pi_{0}(f^{-1}(W^{\prime}))}\omega_{A^{\prime}}^{U}(q_{L^{\prime}})$
and it is enough to prove that $f^{*}(\omega_{W,A})\omega_{A^{\prime}\cap X}^{U}(q_{L^{\prime}})=0$ for all $L^{\prime}\in\pi_{0}(f^{-1}(W^{\prime}))$. We apply $g^{*}$ to obtain
$g^{*}(f^{*}(\omega_{W,A})\omega_{A^{\prime}\cap X}^{U}(q_{L^{\prime}}))=\sum_{L\in\pi_{0}(g^{-1}f^{-1}(W))}\omega_{A}^{V}(q_{L})\sum_{K\in\pi_{0}(g^{-1}(L^{\prime}))}\omega_{A^{\prime}\cap X}^{V}(q_{K})$
but all products $\omega_{A}^{V}(q_{L})\omega_{A^{\prime}\cap X}^{V}(q_{K})$ are zero because $X$ is a dependent set, see Corollary 4.4. This complete the proof. ∎
Lemma 5.7.
Let $X=C\sqcup F$ be a central set of nullity 111 and let $C$ be the unique oriented circuit contained in $X$, $Y$ be a connected component of $\bigcap_{i\in X}H_{i}$, then
@@
We apply Lemma 3.6 to the circuit $C$ in the abelian variety $U=\operatorname{Hom}(\Lambda^{\prime},G)$. Let $p\in Y$ be any point. If $b=1$, we have: